Crew 265 Sol Summary 24 APR 2022

[Crew 265 Sol Summary April 24th]

Crew 265 Sol Summary Report 04/24/22

Sol: 0

Summary Title: Arrival at MDRS

Author’s Name: Marc Levesque, Crew Commander

Mission Status: Nominal

Sol Activity Summary:

Crew 265 arrived at MDRS over the weekend and settled into the Hab this afternoon. Three water runs were made to fill the water tank, with a potential fourth run required before going into sim. Some of the crew tried on their EVA gear, with the rest scheduled for tomorrow. The crew also expressed their appreciation for the fast Internet service.

Look Ahead Plan: The crew will meet with Station Director Rupert to complete their orientation on EVAs and Rovers and will conduct an additional water run if needed. Following that, we expect to go into sim at noon.

Anomalies in work: None

Crew Physical Status: Good and excited to be starting their mission.

EVA: None

Reports to be filed:

Sol Summary

Operations Report

HSO Report

Support Requested: None

Crew 265 Operations Report 24 APR 2022

[Crew 265 Operations Report April 24th]

Crew 265 Operations Report 24-04-2022

SOL: 0

Name of person filing report: Dave Laude

Non-nominal systems: Heating system.

Notes on non-nominal systems: The heating system is now functional after non-functioning with previous crew. I used the first new thermostat that was ordered as that is correct for this system. However, it does not have fan switches like the original did. The main issue is that the exhaust gas fan to the outside cycles on and off because the fan operation pressure sensing switch is stuck closed and so the furnace senses this as a fault. For operation I have to fake the opening and closing during start-up as it is being checked by the electronics. We only use the furnace in the morning for a few hours if very cold in the hab and when I can tend to it. Warm temps are coming soon and so we may not need it. It will not be run overnight as is. No replacement part is planned to be ordered because the entire furnace is to be replaced after this crew season is over.


Spirit rover used: No

Hours: (before EVA)

Beginning charge: (Before EVA)

Ending charge: (On return from EVA, before recharging)

Currently charging: Yes

Opportunity rover used: No


Beginning charge:

Ending charge:

Currently charging: Yes

Curiosity rover used: No


Beginning charge:

Ending charge:

Currently charging: Yes

Perseverance rover used: No


Beginning charge:

Ending charge:

Currently charging: Yes

General notes and comments:

Summary of Hab operations:


Water (static tank): Not full. ~480 gallons. Does this warrant a special trip for water

Water (loft tank): 20 gallons

Water Meter: 0158026.1 units

Static to Loft Pump used – No

Static tank pipe heater (on or off): Off

Static tank heater (On or off): Off

Toilet tank emptied: No

Summary of internet: Awesome!

Summary of suits and radios: Batteries charging. One radio in RAM seems to have a very weak battery battery. Will need to test it. Radio replaced with one from lower hab deck. Testing of some currently unused EVA suit battery capacity in inventory has started. Scott Davis has asked for this on a sample basis to confirm or not that batteries labeled dead are of no use. Will start some radio battery capacity testing soon.

Summary of GreenHab operations: Not used, nothing to report

WATER USE: gallons

Heater: On or Off

Supplemental light: On or Off


Summary of ScienceDome operations: Not used, nothing to report

Dual split: Heat or AC, Off

Summary of RAM operations: Tools used for furnace work. DMMs used for battery checks.

Summary of any observatory issues: Won’t be using, nothing to report

Summary of health and safety issues: Star Link cable has been taped to the floor on the lower deck to prevent tripping upon it.

Questions, concerns and requests to Mission Support: (note trash pickup)

Water (static tank): Not full. ~85% or ~470 gallons. Maybe needs another 6-8". Does this warrant a special trip for water tomorrow or can someone deliver some during the mission?

Minutes ago we smelled natural gas or propane on the lower deck only. The gas detector alarm was not set off. Manual detectors read null and the gas dissipated in about 10 minutes. This same situation has occurred on other crews. We will check periodically for the same occurance.

Crew 245 Final Mission Summary

Crew 245 rotation, planned between April 10th and 23rd, had officially started the simulation on April 12th due to a delay in the delivery of some equipment needed for the experiments and completed the simulation on Friday 23rd morning. During these 10 days of simulation, the crew successfully completed 15 EVAs to perform the targeted experiments and visited several areas in the West, North and East of the habitat. 

The SMOPS crew had initially planned 13 experiments, both passive and active, mainly focused on crew monitoring and support, but two of them were cancelled during the mission for several reasons. The experiments can be divided in four main categories: crew health, space suits and monitoring, space support technology and planetary science. For the first category, the crew had performed a spit test for cortisol measurement to estimate two crew members’ stress levels before and after an EVA in which they had to navigate in an area using only a map (a standard and a drone generated one) and a compass. The saliva samples had been collected by the HSO crew and they will be shipped to a laboratory for the analysis. In the second category, crew gear (prototype flight suit and boots) and wearable sensors have been tested to monitor the movements of crew members during EVAs, data have been downloaded after each EVA and it will be delivered to Mars Planet for the post-processing. During the entire mission, each crew member also wore an undergarment for continuous monitoring of physiological parameters (e.g. heart rate, breathing, oxygen level in the blood). The data have been downloaded by the HSO crew that will post-process it after the mission. The crew has also tested some technologies that will support astronaut activities in future missions, such as purification of air from bacteria that had been installed in the upped deck of the habitat. The air purifier also monitored the quality of the air and the measurement will be compared with the manual measurements performed by the crew Engineer with another device every other day.

In the first days of the mission, a satellite communication system has been installed by the crew (an external antenna next to the RAM and a ground station inside the RAM module) during multiple EVAs, however due to the complexity of the system it was not possible to fully troubleshoot the system and connect with the targeted satellite.

Another tested technology was the 3D scanning of station modules and geological features, data have been collected with the 3D scanner and it will be processed after the mission to generate 3D models of the station and of the geological formations.

The crew also brought a 3D Printer that was used to produce tools in support to other experiments (e.g. a scoop and collection tools for the geological samples) and a drone (a small quadcopter) that was used to map a couple of areas (the station and Tank wash areas). Last but not least, the crew collected and processed twelve geological samples that were processed by the crew scientist that found some magnetic particles and separated them from the main samples. The samples will be shipped to the principal investigator that will assess the origin of the microparticles (if micrometeoroids or simply ferromagnetic material). 

Also ten biological samples were collected and the crew scientist had processed them following the principal  investigator procedure to extract DNA material that will be further analysed by the researcher after the mission.

Throughout the mission we also performed outreach and filming activities: the crew journalist Benjaman is currently working on a documentary on analogue missions and he filmed the crew performing operations with cameras and drones, but also nice feature of the environment around the station, like a sky timelapse. 

Crew 245 Research Report April 22nd

Crew 245 Mid-mission Research Report

– 3D Printed objects utility

Research objectives: The experiment aims to test various 3D-printed objects for EVAs (surface sampling activities, drone’s spare) and medical activities. Each 3D-printed material will be tested immediately after printing, then use during nominal operations. The raw material used for the experiment are from polymers which offer similar mechanical properties as metals. After each use, fill a questionnaire and pictures of the object to analyse the degradation of the object with the use.

Achieved so far: Successfully tested the several geology tools for sample collections and storage. Crew 3D printed some of them before the mission but also during the missions.

– 3D Scan

Research objectives: The experiment aims to perform a detailed scan of the Habitat (internal and external) and generate a detailed 3D map of it. Potentially, other geological features might be scanned and mapped.

Achieved so far: Crew performed a scan of the internal part of the habitat, of the RAM and Science dome together with four geological formations during EVAs. The crew has downloaded the data and advanced in the post-processing of the data with a dedicated software and a python script adapted by the crew.

– Air Purification for Life Support systems

Research objectives:
The experiment focuses on three main parts or “sub-experiments”:
• Experiment 1: monitoring the quality of the air by measuring the degree of contamination (or particulate matter PM) carried by the crew members.
• Experiment 2: Bacteriological detector for the evaluation of the bacterial concentration.
• Experiment 3: Bacterial cultivation in a controlled environment.

Achieved so far: the air purifier has been installed in the crew quarters of the habitat and it was activated right after. In addition to the daily filtering routine, the crew performs every other day manual measurements with a separate device to compare the two measurements. The crew has also performed the last part of the experiment (#3) in a controlled environment

– Biomonitors for continuous care

Research objectives:
The purpose of the experiment is to continuously monitor the health of the crew members through a biomonitor device that includes integrated sensors designed to measure the body activity, heart rate, and body temperature and pressure of the crew members.
The members wear the biomonitor sensors undersuite and as a head band.
The data collected from the body monitoring are processed through a dedicated available software, and then transmitted to the principal investigators to be analyzed.

Achieved so far: Crew wore the sensors (on a T-shirt and headband) for the whole mission with the Crew HSO checking the correct functioning and recording of the data and values from time to time. The downlinked data will be processed post mission.

– BGM Mark 1 suits

Research objectives: Tests of a flightsuit simulator designed by Radici and shoes provided by calzaturificio King. These parts are worn and tested with sensors to measure quantities relevant for spacesuit simulator R&D. In particular, crew members wear an Emotibit sensor on their arm, and a custom made accelerometer on the right foot. A helmet with integrated air filtering and ventilation system is tested.
This experiment is part of a larger program for the development of a space suit simulator for analogue missions.

Achieved so far: crew members have been wearing the experimental flight suits and boots during EVAs. The helmet has been also successfully tested. Crew continues to wear the sensors (arm and shoes) during EVAs and the Crew Executive Officer (XO) downloads the data at the end of each EVA session. The crew will continue to wear them for the rest of the mission and data will be processed post mission. Feedback on the use of suit, boots and helmet are being collected by the XO.

– Evaluation of drone operations and questionnaires

Research objectives:
Mapping (1)* of terrain and inspection (3) of the station and its elements (e.g. solar arrays) during Extra-Vehicular Activities (EVAs), also simulating possible search and rescue (2) events.
Acquisition of the martian base camp outside environment monitoring sandstorms through a warning weathercast system (4).
For the scenarios, a commercial drone is used for photogrammetry and surface analysis and exploration. The images produced are used in support of other experiments.
Note: () Numbers indicate priority

Achieved so far: Crew performed a test for the mapping function and a first full mapping session on Sol7. The generated map was used for the navigation test required for the Spit test experiment.

– Micrometeorites from the Solar System

Research objetives: The main purpose of the experiment is find the extra-terrestrial materials on the earth’s surface through the search of micrometeorites in the analog field.
The crew will collect minimum 10 and maximum 30 geological samples in different locations in the area of the Mars Research Desert Station that will be then analysed inside the station to search for micrometeorites.
All the samples will be stored and delivered to the PI after the mission.
Similar investigation and sample collection will be carried out in other analog environments (e.g. AMADEE20 performed by the Austrian Space Forum) and these different analog areas will be compared by extra-terrestrial traces. The final goal is to establish an analog mission micrometeorite archive.

Achieved so far: Crew collected a total of twelve samples out of the ten minimum required. The collected samples have been analysed by the crew scientist that has found several particles that were attracted by the magnetic field generated by the portable magnet. Those particles have been separated from the main samples and they will be further analysed by the principal investigator.

– Portable DNA analyser in search for Extraterrestrial life

Research objetives: This experiment aims to optimize research methods based on the restrictions of human-robotic exploration mission. Here, we focus on the easy-to-do methods including classical microbiology and NGS (next generation sequencing) techniques to identify extremophile bacteria, archaea, tardigrades, and slimemold. During this mission we aim to analyse sand samples on traces of microbial DNA. The experiment contains the following steps (1) sample collection; (2) DNA extraction; (3) sequencing, (4) analyses and interpretation (post-mission)

Achieved so far: Crew collected the maximum required number of samples (ten) and they were processed by the crew Scientist that has extracted the DNA by following the procedure prepared by the principal investigator. Unfortunately, due to an anomaly with one of the components of the portable sequencer, the crew was not able to sequence the DNA on-site, but the library of extracted DNA samples has been shipped to the principal investigator that will perform the troubleshooting and complete the DNA analysis.

– Satellite communication and data exchange

Research objetives: The purpose of the experiment is measuring the communications level between the GCS and the on-orbit satellite (D-orbit) through the amateur radio frequency. The software tool required by the experiment is also provided by D-Orbit.

Achieved so far: The crew installed the antenna and the ground station, it also attempted to connect a few times with the satellite during its pass above the station but without success. Troubleshooting activities have been attempted, but without much luck.

– Spit cortisol test

Research objectives: Measuring the cortisol level affecting the crew members. This test plays an important role in understanding (but also preventing) potential health risks, such as high blood pressure, high blood sugar, muscle weakness etc. The test could be done through blood samples combined with cyclomatic tests for the crew.

Achieved so far: The collection of the samples for the cortisol tests have been performed on the two targeted subjects after the successful execution of the navigation tests using a standard and a drone generated maps.

– EEG set during meditation

Research objectives: The purpose of this experiment consists of theoretical and practical study of the use of portable EEG headsets for neurofeedback experiments in isolated, confined and extreme environments. The aim of the research is to inquire about new non-intrusive methodology to countermeasure stress in isolation to enhance human performances in future long length space missions.

Achieved so far: due to technical issues with the EEG sent and a mishap in communication with another partner, the experiment was canceled.

Crew 245 Operations Report April 22nd

category operations-report]

Crew 245 Operations Report 22-04-2022
SOL: 12
Name of the person filing report: Luca Rossettini

Non-nominal systems: Heating system.
Notes on non-nominal systems: –
Spirit rover used: No
Hours: 173
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes
Opportunity rover used: No
Hours: 95.6
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes
Curiosity rover used: No
Hours: 185.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes
Perseverance rover used: No
Hours: 235.6
Beginning charge: 100%
Ending charge: -%
Currently charging: –
General notes and comments: Not used today
Summary of Hab operations:
Water (static tank): 323 gallons
Water (loft tank): 12 gallons
Water Meter: 157955.8
Static to Loft Pump used: Yes
Static tank pipe heater: Off
Static tank heater: Off
Toilet tank emptied: No
Summary of internet: nothing to report
Summary of suits and radios:
Summary of GreenHab operations:
WATER USE: 0 gallons
Heater: Off
Supplemental light: Off
Summary of ScienceDome operations: Nothing to report
Dual split (Heat/AC): Off
Summary of RAM operations: nothing to report
Summary of any observatory issues: Not used, nothing to report
Summary of health and safety issues: nothing to report
Questions, concerns and requests to Mission Support: –

Supplemental Operations Report April 24th

Supplemental Operations Report 24 April 2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: n/a

Generator: Still being manually run at night from about 9 pm to 7:30 am

ScienceDome Dual Split: Off

Solar—Charge controllers off at night.

Solar— Nominal

Notes on power system: Nothing to report

Propane Reading, station tank – 57 %

Propane Reading, director tank— 58 %

Propane Reading, intern tank— 68 %

Propane Reading, generator—43 %

Ethanol Free Gasoline – 0 gallons.

Water (Outpost tank) – 500 gallons

Sojourner rover used: no

Hours: 167.7

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: All rovers were serviced. The batteries were cleaned, neutralized and the cells of the batteries were refilled.

Spirit showed some corrosion and batteries were evenly depleted of water.

Curiosity showed some corrosion and batteries were unevenly depleted of water. The driver’s side had water levels equal to that of Spirit, but the passenger’s side were much more depleted. I think they were accidentally not filled the last time.

Percy showed limited corrosion and batteries were evenly depleted of water. One battery took some effort to refill due to a rusted bolt.

Opportunity showed some corrosion and batteries were evenly depleted of water.

Sojourner showed heavier corrosion and batteries were evenly depleted of water.

It took about 70 gallons of water to clean them and 13 gallons of distilled water to fill them. They will need to be serviced again during the work party.

ATV’s Used: (Honda, 350.1, 350.2, 300): none

Reason for use: n/a

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0 hours

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town for garbage and water

CrewCar used and why, where? Yes, to Grand Junction. I fixed the broken piece on the interior driver’s side by the windshield. It just needed to pushed securely back into place.

Luna used and why, where? Yes, to Denver. Got a warning that the tire sensor is inoperable and needs to be reset. I will work on figuring this one out. Also the tires showed four different pressures right before this, so I will adjust them as needed.

Campus wide inspection, if action taken, what and why? Nothing to report

General notes and comments: Dave and Sergii got here on Saturday and helped me do things before the arrival of the crew. This was deeply appreciated.

Summary of general operations: Nothing to report

Summary of internet: We removed the multiple sets of HudgesNet and old internet wiring throughout the Hab. We also removed the blue ethernet wire that had been put in place for the webcams that could not be used. Things look much neater as a result.

Summary of suits and radios: Dave is working with Scott on some troubleshooting of some of the suits and batteries.

Summary of Hab operations: Dave did some troubleshooting of the furnace and it is semi-operational. It can be turned on to warm the Hab but it is not completely safe as he had to bypass an exhaust sensor to do so. In addition, the new thermostat does not have an on/off switch and turns on at 50 degrees, so it will need to be disabled after each use by disconnecting the wires. This is a good enough work around for this crew, and as they are the last crew that will need it before a new furnace is installed, it will work. They will not run the furnace at night and in a few days it should be warm enough to retire the furnace forever. Dave also reset the water alarm in the loft tank.

Summary of Outpost operations: The new pump and hose have arrived for the water tank and I will set it up when I need more water.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: I dropped a 10mm wrench from the newer set between the batteries of Percy and could not retrieve it. There is an older 10 mm single wrench that can replace it until we can get it out from where it is wedged.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report


Copyright © The Mars Society. All rights reserved. | Main Site